Suppose the production function for a firm is given by: q=2L +4K. If the firm currently has 20 units of capital (K) and 10 units of labor (L), then calculate the Marginal Rate of Technical Substitution (MRTS). (Round to the nearest two decimal places if necessary.) Answer:
Answer #1


Production function is as follows -

q = 2L + 4K

Marginal rate of technical substitution between two inputs is the ratio of their respective marginal products.

So,

Calculate the marginal product of labor (MPL) -

MPL = dq/dL = d(2L+4K)/dL = 2

Calculate the marginal product of capital (MPK) -

MPK = dq/dK = d(2L+4K)/dK = 4

Calculate MRTSLK -

MRTSLK = MPL/MPK

MRTSLK = 2/4 = 0.50

Thus,

The Marginal Rate of Technical Substitution (MRTSLK) is 0.50.